A bank loaned out $67,000, part of it at the rate of 12% per year and the rest at a rate of 6% per year. If the interest received was $6060, how much was loaned at 12%?
Answer by richwmiller(17219)

x+y=67000,
0.12*x+0.06*y=6060
x=67000-y
0.12*(67000-y)+0.06*y=6060
8040-0.12y+0.06*y=6060
-0.06*y=-1980
y=33000 at 6%
x=67000-y
x=34000 at 12%
check
0.12*34000+0.06*33000=6060
4080+1980=6060
6060=6060
ok